# AdaCAD: Adaptively Decoding to Balance Conflicts between Contextual and Parametric Knowledge

### v1

@@ 4.2 Main Results
In Table 2, we investigate how AdaCAD can improve performance on longer-form generation, showing results on three summarization tasks, CNN-DM, Xsum, and TofuEval. For TofuEval, AdaCAD demonstrates substantial improvements, particularly excelling in marginal topics (i.e. topics not central to the document) where it outperforms greedy decoding, CAD, and COIECD by 14.60, 2.21, and 17.48 points in terms of AlignScore – a measure of faithfulness – respectively. This highlights AdaCAD’s ability to handle diverse topics and maintain factual consistency, especially when prompted to focus on a marginal topic; qualitatively, we see in Fig. 4 that these improvements are driven by less hallucination on the part of AdaCAD.
@@
On CNN-DM, AdaCAD achieves the highest ROUGE-L score of 25.42, surpassing greedy decoding, CAD, and COIECD by 0.49, 0.66, and 1.95, respectively. In terms of factual consistency, AdaCAD also leads with an AlignScore of 94.97. On XSum, AdaCAD also outperforms all baselines across all metrics. For instance, AdaCAD achieves an average improvement of 1.38, and 4.67 points in BERT-P, and AlignScore, respectively. For BERT-P metric on CNN-DM, AdaCAD outperforms CAD and COIECD and is slightly lower than Greedy decoding; as mentioned in Section 4.1, this may be a result of a lack of conflict in these datasets, which are at least partly included in large pretraining corpora. These improvements indicate that AdaCAD’s dynamic adjustment mechanism is effective in long-form generation, allowing it to balance context and parametric knowledge.
@@
5 Analysis
Setup.
In Table 1, we find that CAD underperforms AdaCAD, COIECD, as well as greedy decoding on most QA datasets, except NQ-Swap, wherein by design there is a high degree of conflict between the context and LLM’s parametric knowledge Longpre et al. (2021). We hypothesize that on more realistic datasets, the trailing performance of CAD stems from its inability to account for instances with low or minimal conflict. To test this hypothesis, we evaluate all methods on examples designed to have minimal conflict, i.e., where the model’s internal representation aligns well with the context. Specifically, we generate a dataset of synthetic non-conflicting data called NQ-Synth: we sample 500 questions from Natural Questions and then prompt the Llama-3-70B to generate the answer for each question. We replace the gold answer entity in the context with the generated answer by regex, thus, making the context consistent with the LLM’s internal knowledge. Finally, we evaluate Llama-3-70B on NQ-Synth (which, by design, has no conflict) and on NQ-Swap (which has conflict in every example). See Table 5 in Appendix B for examples of NQ-Swap and NQ-Synth.
@@ 5.1 Performance comparison on instances with higher and lower degrees of conflict
The results are shown in Table 3. Consistent with our hypothesis, we observe that in the absence of conflict (i.e., on NQ-Synth), CAD substantially degrades performance by $\approx\!24\%$ relative to greedy decoding, while AdaCAD maintains a comparable performance. Although COIECD seeks to detect conflict and operates in two distinct decoding modes for high and low confllict, it also underperforms in non-conflict scenarios, falling $2.8\%$ behind AdaCAD. However, in cases of high conflict (i.e., NQ-Swap), where greedy decoding struggles and yields dramatically lower accuracy,  AdaCAD improves over greedy decoding by $11.2\%$, while COIECD cannot handle high-conflict examples as well, lagging behind AdaCAD by $12\%$. To further investigate how AdaCAD balances instances with lower and higher degrees of conflict, we compute $\alpha^{\textsc{jsd}}_{\mathrm{max}}$, which is the maximum $\alpha^{\textsc{jsd}}_{t}$ value across tokens, for both datasets. Indeed, we find that $\alpha^{\textsc{jsd}}_{\mathrm{max}}$ adapts to the amount of conflict, with an average value of $0.45$ on NQ-Swap with a higher level of conflict, and substantially lower value ($\alpha^{\textsc{jsd}}_{\mathrm{max}}\!=\!0.28$) on NQ-Synth which does not contain any conflict by design.
@@ 5.2 PMI does not adequately address conflict
As described in Sections 3.2 and 3.3, both CAD and AdaCAD compute the PMI $\left(p_{\theta}(y_{t}\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})\big{/}p_{\theta}(y_{t}\mid\boldsymbol{x},\boldsymbol{y}_{<t})\right)$ between the LLM’s output distributions with and without external context $\boldsymbol{c}$. However, CAD relies solely on the PMI term to balance the level of conflict, whereas in AdaCAD, we compute $(\text{PMI})^{\alpha_{t}^{\textsc{jsd}}}$ where both PMI and $\alpha_{t}^{\textsc{jsd}}$ adapt with the degree of conflict. In cases of low conflict, the LLM’s output distributions should in principle be indistinguishable with and without the context, rendering $\text{PMI}\!\approx\!1$, i.e., resorting to greedy decoding irrespective of the value of $\alpha$ (cf. Eq. 1). However, in practice, we find that, even in cases of minimal conflict, the PMI term reranks the tokens in the head of the LLM’s distribution, resulting in poor performance for CAD.
Setup.
To test how the PMI term accounts for conflict, we measure the amount of reranking (among tokens) done by CAD and AdaCAD relative to the greedy distribution $p_{\theta}\left(y_{t}\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t}\right)$. We compute the Spearman rank-order correlation coefficient $\rho$ between the greedy distribution and output distribution from CAD and AdaCAD (with scaling factors PMI and (PMI)${}^{\alpha_{t}^{\textsc{jsd}}}$ respectively). We restrict the measurement to the top-20 tokens (averaged across decoding steps) on NQ-Swap and NQ-Synth. [^6] Since the rank of low-probability tokens does not influence the generation, we focus on the top-20 tokens with the highest probability at each step. A lower rank coefficient means more tokens have been reranked by the adjustment method, while a higher coefficient means fewer tokens have been reranked. Intuitively, a method sensitive to the degree of conflict should yield a lower rank correlation (more perturbation) when the amount of conflict is high (i.e., on NQ-Swap), and higher rank correlation (less perturbation) in cases of low conflict (i.e., on NQ-Synth). We also compute the absolute difference or sensitivity, $\lvert\Delta\boldsymbol{\rho}\rvert$ between the two $\boldsymbol{\rho}$ values of NQ-Swap and NQ-Synth. A larger $\lvert\Delta\boldsymbol{\rho}\rvert$ indicates that the method is more effective at distinguishing between conflicting and non-conflicting data, i.e., more sensitive to the degree of conflict in instances.
Results in Table 4 demonstrate that CAD, which only relies on the PMI term to offset conflicts, perturbs the greedy distribution to roughly the same extent in the presence or absence of conflict, with similar correlation coefficients on NQ-Swap and NQ-Synth. This minimal difference $\lvert\Delta\boldsymbol{\rho}\rvert$ suggests that CAD applies a similar level of adjustment regardless of the amount of conflict, leading to over-correction for non-conflicting examples. On the other hand, the correlation coefficient of AdaCAD is higher on NQ-Synth than on NQ-Swap (0.94 vs. 0.86), indicating more perturbation to the greedy distribution in the presence of conflict. Additionally, the sensitivity to conflict ($\lvert\Delta\boldsymbol{\rho}\rvert$) of AdaCAD is substantially larger (8×8$\times$8 ×) than that of CAD, highlighting AdaCAD’s superior ability to distinguish between conflicting and non-conflicting examples. Note that AdaCAD has a higher $\rho$ in both settings, indicating that it perturbs the LLM’s distribution less overall.
@@
Figure 3: Qualitative example of different methods on NQ. Text about correct answers is highlighted in blue, while text about unrelated facts and wrong answers is highlighted in red. AdaCAD produces the correct answer, while CAD generates unrelated outputs and COIECD fails to detect the conflict and generates the same incorrect answer as greedy decoding.
Figure 4: Qualitative example of different methods on TofuEval. Text unsupported by the transcripts is highlighted in red, while consistent and relevant text is highlighted in blue. AdaCAD generates a more faithful summary whereas other methods tend to hallucinate details.
@@ 6 Discussion and Conclusion
In naturalistic scenarios with datasets containing a heterogeneous mix of examples with and without conflicts between parametric and contextual knowledge, existing decoding methods, including CAD, fail to adapt to changing amounts of conflict and in fact can lead to reduced performance. Although larger and more performant models can store more information in their parametric knowledge – thus leading to less and less conflict as models improve – there will still always be gaps between the model and the actual state of the world (e.g., because of time cutoffs). This means that models will encounter both low- and high-conflict scenarios, no matter their strength.
To this end, we introduce AdaCAD, a simple yet effective dynamic decoding method that utilizes Jensen-Shannon divergence to dynamically model the degree of conflict for a given example (and a given timestep) and automatically balance the contrast between contextual and parametric knowledge. On diverse QA datasets, we show that AdaCAD combines the advantages of greedy decoding and context-aware decoding, resulting in improved performance. Additionally, experiments on summarization datasets demonstrate that AdaCAD enhances both the quality and factuality of the generated summaries, while other methods tend to hallucinate details. Analysis reveals that AdaCAD effectively mitigates the overcorrection observed in CAD by dynamically adjusting the weight of contextual knowledge based on the degree of conflict. Lastly, AdaCAD consistently outperforms other hybrid decoding strategies that detect conflict, COIECD, demonstrating superior flexibility and accuracy across varying levels of conflict.
@@ Acknowledgements
We would like to thank David Wan for feedback on our summarization experiments. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, and NSF-CAREER Award 1846185. The views contained in this article are those of the authors and not of the funding agency.

### v2

@@ 4.2 Main Results
In Table 2, we investigate how AdaCAD can improve performance on longer-form generation, showing results on three summarization tasks, CNN-DM, Xsum, and TofuEval. For TofuEval, AdaCAD demonstrates substantial improvements, particularly excelling in marginal topics (i.e., topics not central to the document) where it outperforms greedy decoding, CAD, COIECD, and ConfCD by 14.60, 2.21, 17.48, and 2.56 points in terms of AlignScore – a measure of faithfulness – respectively. This highlights AdaCAD’s ability to handle diverse topics and maintain factual consistency, especially when prompted to focus on a marginal topic; qualitatively, we see in Fig. 4 that these improvements are driven by less hallucination on the part of AdaCAD.
@@
On CNN-DM, AdaCAD achieves the highest ROUGE-L score of 25.42, surpassing greedy decoding, CAD, COIECD, ConfCD by 0.49, 0.66, 1.95, and 1.48, respectively. In terms of factual consistency, AdaCAD also leads with an AlignScore of 94.97. On XSum, AdaCAD also outperforms all baselines across all metrics. For instance, AdaCAD achieves an average improvement of 1.43, and 5.46 points in BERT-P, and AlignScore, respectively. For BERT-P metric on CNN-DM, AdaCAD outperforms all contrastive decoding baselines and is slightly lower than Greedy decoding; as mentioned in Section 4.1, this may be a result of a lack of conflict in these datasets, which are at least partly included in large pretraining corpora. These improvements indicate that AdaCAD’s dynamic adjustment mechanism is effective in long-form generation, allowing it to balance context and parametric knowledge.
5 Analysis
Setup
In Table 1, we find that CAD underperforms AdaCAD, COIECD, as well as greedy decoding on most QA datasets, except NQ-Swap, wherein every instance by design has a high degree of conflict Longpre et al. (2021). We hypothesize that on more realistic datasets, the trailing performance of CAD stems from its inability to account for instances with low or minimal conflict. To test this hypothesis, we evaluate all methods on examples designed to have minimal conflict, i.e., where the model’s internal representation aligns well with the context. Specifically, we generate a dataset of synthetic non-conflicting data called NQ-Synth: we sample 500 questions from Natural Questions and then prompt the Llama-3-70B to generate the answer for each question. We replace the gold answer entity in the context with the generated answer by regex, thus, making the context consistent with the LLM’s internal knowledge. Finally, we evaluate Llama-3-70B on NQ-Synth and on NQ-Swap. See Table 6 in Appendix B for examples of NQ-Swap and NQ-Synth.
@@ 5.1 Performance comparison on instances with higher and lower degrees of conflict
Consistent with our hypothesis, in Table 3, we observe that in the absence of conflict (on NQ-Synth), CAD substantially degrades performance by $\approx\!24\%$ relative to greedy decoding, while AdaCAD maintains a comparable performance. Although COIECD seeks to detect conflict and operates in two distinct decoding modes for high and low confllict, it also underperforms in non-conflict scenarios, falling $2.8\%$ behind AdaCAD. However, in cases of high conflict (NQ-Swap), where greedy decoding yields dramatically lower accuracy,  AdaCAD improves over greedy decoding by $11.2\%$, while COIECD cannot handle high-conflict examples as well, lagging behind AdaCAD by $12\%$. To further investigate how AdaCAD balances instances with lower and higher degrees of conflict, we compute $\alpha^{\textsc{jsd}}_{\mathrm{max}}$, which is the maximum $\alpha^{\textsc{jsd}}_{t}$ value across tokens, for both datasets. Indeed, we find that $\alpha^{\textsc{jsd}}_{\mathrm{max}}$ adapts to the amount of conflict, with an average value of $0.45$ on NQ-Swap with a higher level of conflict, and substantially lower value ($\alpha^{\textsc{jsd}}_{\mathrm{max}}\!=\!0.28$) on NQ-Synth which does not contain any conflict by design.
@@ 5.2 PMI does not adequately address conflict
As described in Sections 3 and 3, both CAD and AdaCAD compute the PMI between the LLM’s output distributions with and without external context $\boldsymbol{c}$. However, CAD relies solely on the PMI term to balance the level of conflict, whereas in AdaCAD, we compute $(\text{PMI})^{\alpha_{t}^{\textsc{jsd}}}$ where both PMI and $\alpha_{t}^{\textsc{jsd}}$ adapt with the degree of conflict. In cases of low conflict, the LLM’s distributions should in principle be the same with and without context, rendering $\text{PMI}\!\approx\!1$, i.e., resorting to greedy decoding for any value of $\alpha$ (cf. Eq. 1). However, in practice, we find that, even with minimal conflict, the PMI term reranks the tokens in the head of the LLM’s distribution, resulting in poor performance for CAD.
@@
Setup
To test how well the PMI term accounts for conflict, we measure the amount of reranking (among tokens) done by CAD and AdaCAD relative to the greedy distribution. We compute the Spearman rank-order correlation coefficient $\rho$ between the greedy distribution and output distribution from CAD and AdaCAD (with scaling factors PMI and (PMI)${}^{\alpha_{t}^{\textsc{jsd}}}$ respectively). We restrict the measurement to the top-20 tokens (averaged across decoding steps) on NQ-Swap and NQ-Synth. [^6] As the rank of low-probability tokens does not influence the generation, we focus on the top-20 tokens at each step. Intuitively, a method sensitive to the degree of conflict should yield a lower rank correlation (more perturbation) when the amount of conflict is high (on NQ-Swap), and higher rank correlation (less perturbation) in cases of low conflict (on NQ-Synth). To this end, we compute the absolute difference or sensitivity, $\lvert\Delta\boldsymbol{\rho}\rvert$ between the two $\boldsymbol{\rho}$ values of NQ-Swap and NQ-Synth. A larger $\lvert\Delta\boldsymbol{\rho}\rvert$ indicates that the method is more effective at distinguishing between conflicting and non-conflicting data, i.e., more sensitive to the degree of conflict in instances.
Datasets CAD (tuned $\alpha$) AdaCAD
NQ 44.35 (0.25) 45.47
TriviaQA 79.60 (0.25) 82.50
PopQA 78.19 (0.25) 81.34
HotpotQA 46.81 (0.50) 50.53
TabMWP 46.90 (0.50) 53.00
Average 59.17 62.57
Table 5: Performance of CAD with tuned $\alpha$ and AdaCAD on QA datasets with Llama3-8B.
Figure 3: Qualitative example from NQ. Informative text is highlighted in blue, while text about unrelated facts and wrong answers is highlighted in red. AdaCAD produces the correct answer, while CAD generates unrelated outputs and COIECD fails to detect the conflict and generates the same incorrect answer as greedy decoding.
Figure 4: TofuEval: Text unsupported by the transcripts is highlighted in red, while consistent and relevant text is highlighted in blue. AdaCAD generates a faithful summary whereas other methods tend to hallucinate details.
Results in Table 4 demonstrate that CAD, which only relies on the PMI term to offset conflicts, perturbs the greedy distribution to roughly the same extent ($\rho$) in the presence or absence of conflict, i.e., on NQ-Swap and NQ-Synth, respectively. This minimal difference in $\lvert\Delta\boldsymbol{\rho}\rvert$ suggests that CAD is agnostic to the amount of conflict, leading to over-correction for non-conflicting examples. On the other hand, the correlation coefficient of AdaCAD is higher on NQ-Synth than on NQ-Swap (0.94 vs. 0.86), indicating more perturbation to the greedy distribution in the presence of conflict. Additionally, the sensitivity to conflict ($\lvert\Delta\boldsymbol{\rho}\rvert$) of AdaCAD is substantially larger (8×8$\times$8 ×) than that of CAD, highlighting AdaCAD’s superior ability to distinguish between conflicting and non-conflicting examples. Note that AdaCAD has a higher $\rho$ in both settings, indicating that overall, it perturbs the LLM’s distribution to a lesser extent.
5.3 Tuning $\alpha$ of CAD for each dataset
Since AdaCAD does not require validation data to tune the value of $\alpha$, we set CAD’s $\alpha=1$ (tuned on NQ-SWAP) for QA datasets following Shi et al. (2024), which may explain the strong performance of CAD on NQ-SWAP and low performance on other datasets. To further underscore the advantages of AdaCAD over CAD, we compare AdaCAD (untuned) to a CAD baseline with a tuned $\alpha$ value. Specifically, we tune CAD’s $\alpha$ using a validation set of 500 instances (randomly sampled from the train set) for each dataset.
Table 5 shows that AdaCAD achieves an average improvement of $3.4\%$ (absolute) over CAD even when $\alpha$ is tuned. We hypothesize that AdaCAD’s superior performance stems from varying the level of adjustment adaptively depending on the underlying instance, whereas a tuned-$\alpha$ CAD still uses the same $\alpha$ uniformly for all instances and does not adjust according to varying degrees of conflict among instances. Moreover, while tuning CAD’s $\alpha$ for each dataset might improve performance in a controlled setting, such tuning does not scale well to real-world scenarios wherein models encounter a mix of user queries – some with high conflict and others with low or no conflict – and this categorization is not known a priori.
@@ 6 Discussion and Conclusion
In naturalistic scenarios with mixed datasets containing examples with and without knowledge conflicts, existing decoding methods, including CAD, fail to adapt to changing amounts of conflict and in fact can lead to reduced performance. Although larger and more performant models can store more information in their parametric knowledge – thus leading to less and less conflict as models improve – there will still always be gaps between the model and the actual state of the world (e.g., because of time cutoffs). This means that models will encounter both low- and high-conflict scenarios, no matter their strength.
To this end, we introduce AdaCAD, a simple yet effective dynamic decoding method that uses Jensen-Shannon divergence to dynamically model the degree of conflict for a given example (and timestep) and automatically balance the contrast between contextual and parametric knowledge. On diverse QA datasets, we show that AdaCAD combines the best of greedy decoding and context-aware decoding, improving performance. Additionally, experiments on summarization demonstrate that AdaCAD enhances both the quality and factuality of generated text, while other methods tend to hallucinate details. Lastly, AdaCAD consistently outperforms COIECD, another hybrid decoding strategies that detects conflict. Our analysis reveals that AdaCAD mitigates the overcorrection seen in CAD by dynamically adjusting the weight of contextual knowledge based on the degree of conflict.
Limitations
Since our proposed method AdaCAD is based on CAD, it requires access to output logits from LLMs to calculate the difference between output probabilities with and without context. However, API-based LLMs like GPT-4 often do not provide output logits, making it challenging to directly apply logit-based methods like AdaCAD and CAD to fully black-box models. Additionally, our experiments focus on English datasets and pre-trained models; as LLMs become available for other languages, future research will be needed to explore the interactions between language and knowledge conflict. We do not foresee any particular risks associated with the application of our method.
@@ Acknowledgements
We would like to thank David Wan for feedback on our summarization experiments and the anonymous reviewers for their feedback. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, and NSF-CAREER Award 1846185. The views contained in this article are those of the authors and not of the funding agency.
@@ B.2 Summarization Datasets
B.3 Licenses
Datasets are released under the following licenses:
  • Natural Questions: Apache-2.0 license
  • NQ-Swap: MIT license
  • TriviaQA: Apache-2.0 license
  • PopQA: MIT license
  • HotPotQA: Apache-2.0 license
  • TabMWP: MIT license
  • CNN-DM: Apache-2.0 license
  • XSum: MIT license
  • TofuEval: MIT license
The models we use have the following licenses: